# Karma: An Experimental Study

### v1

@@ Abstract.
A system of non-tradable credits that flow between individuals like karma, hence proposed under that name, is a mechanism for repeated resource allocation that comes with attractive efficiency and fairness properties, in theory. In this study, we test karma in an online experiment in which human subjects repeatedly compete for a resource with time-varying and stochastic individual preferences or urgency to acquire the resource. We confirm that karma has significant and sustained welfare benefits even in a population with no prior training. We identify mechanism usage in contexts with sporadic high urgency, more so than with frequent moderate urgency, and implemented as an easy (binary) karma bidding scheme as particularly effective for welfare improvements: relatively larger aggregate efficiency gains are realized that are (almost) Pareto superior. These findings provide guidance for further testing and for future implementation plans of such mechanisms in the real world.
conference: ; ;
1Automatic Control Lab, ETH Zurich. Emails: {elokdae,bsaverio,dorfler}@ethz.ch
2Department of Sociology, University of Zurich. Email: heinrich.nax@uzh.ch
Research supported by NCCR Automation, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number $180545$).
1. Introduction
Efficiency and fairness in determining who gets what and when are the two major objectives of economists thinking about scarcity, and many interesting matchmaking and market design solutions have been proposed (Roth, 2015). An important class of allocation problems is when goods are repeatedly allocated in the population. In situations like these, it is often the case that one person relative to another gets higher utility from the good in one period but lower utility in another (which we shall capture by the notion of urgency). A recent innovation to address exactly those kinds of allocation problems goes under the name karma (Vishnumurthy et al., 2003; Elokda et al., 2023; Vuppalapati et al., 2023). The karma mechanism mirrors what Western popular culture associates with the notions of karma and samsara stemming from Indian religions according to which one’s deeds in the present (karma) affect the quality of one’s future life (phala) and thereafter (samsara) (Reichenbach, 1990; Kyabgon, 2015). The mechanism is implemented via individual accounts of non-tradable credits called karma. Individuals may bid some amount from their current account at every instance of resource allocation. The highest bidders win, obtain priority for the resource, pay their bids, which are then transferred to the accounts of those who did not win.
Such mechanisms have been proposed previously in non-human contexts to manage scarcity in, for example, traffic (Elokda et al., 2022), transportation (Censi et al., 2019; Salazar et al., 2021), cloud computing (Vuppalapati et al., 2023), and peer-to-peer networks (Vishnumurthy et al., 2003). In these works, agents are assumed to follow perfect Nash equilibrium as they can be programmed to do so. In human contexts, the most well-known example of a karma-like mechanism is the ‘choice system’ adopted by U.S. food banks (Prendergast, 2022), where the role of karma tokens is taken by ‘shares.’ As with karma, food banks under the choice system bid shares to obtain priority for available food donations, and the total payment of shares is redistributed at the end of each day. This real-world implementation of a karma-like system indeed resulted in significant aggregate gains in terms of food bank participation and fluidity of donations. In this paper, we move to a controlled test of such a mechanism as the behavioral uptake by the human actors involved as well as the distributional consequences continue to remain unstudied experimentally.
Karma and the aforementioned choice system are different from other token-based mechanisms that have been implemented and proposed in the past, including those that use tradable credits (Cramton and Kerr, 2002; Sovacool, 2011; Yang and Wang, 2011), non-tradable vouchers (Daniels and Trebilcock, 2013; Hsieh and Urquiola, 2006; Kim et al., 2021), or so-called artificial currencies or scrips (Johnson et al., 2014; Gorokh et al., 2021a, b; Banerjee et al., 2023). In these other mechanisms, some type of voucher or token is typically issued periodically, e.g., at the beginning of every week or month, to be consumed over the period. Karma offers three distinguishing features in comparison to these mechanisms. First, karma is non-tradable even in a black market, as it is associated to individual accounts and can be transferred only in fixed resource allocation instances. Second, one karma token does not equal one resource unit; the price instead arises dynamically from bidding. This has the dual benefit of making the system indifferent to the exact amount of karma in it, and enabling individuals to express their time-varying private urgency. Third, karma forms a closed economy: it is transferred among individuals in a manner that preserves the karma per capita over time. This is advantageous for both system and individuals. From the system perspective, the karma mechanism can be infinitely repeated with minimal central intervention, as no central endowment of karma is needed other than the initial endowment. From the individual perspective, karma is forgiving as it enables individuals to immediately replenish their accounts through yielding, rather than waiting a potentially very long time for the next central endowment.
@@ 1. Introduction
Figure 1 illustrates how karma works out to the benefit of everyone at hand of an example involving three intersection encounters. Let’s start with the encounter in the top center. The high-urgency lila car has a current karma account of 9 and bids 4, thus outbidding the low-urgency blue car whose karma account is also 9 but bids 2. As a result, the blue car’s karma account goes up by 4 to 13. Let’s move along clockwise. Blue now happens to have high urgency and bids 4, thus outbidding and getting priority over orange whose karma account is nine and bids 3. Now the orange karma goes up by 4 to 13. Let’s move along clockwise. Orange now has high urgency and bids 4, thus outbidding low-urgency lila who has 5 karma left and bids 1. Thus the circle closes, et cetera, et cetera.
Figure 1. Dynamic resource allocation with karma.
The example illustrates that karma is similar in spirit to the ‘universal currency’ of reputation (Milinski, 2016), but makes the direct linkages and transfers between users’ karma accounts explicit, while reputation is something that tends to exist in humans’ individual perceptions of one another informally. Similar to how ‘image scoring,’ which is an example of an explicit reputation mechanism, formalizes the notion of indirect reciprocity in reputation-based interactions (Nowak and Sigmund, 1998), karma can be viewed as the formal equivalent of ‘fair turn-taking’ in repeated resource allocation tasks. While image scoring has been subjected to the behavioral test in the past (Wedekind and Milinski, 2000), it remains to do so for karma.
Further, the example also illustrates another feature of karma, as was mentioned above, that distinguishes karma from fixed-value tokens and from other forms of rules ensuring turn-taking: it allows individuals to dynamically express how intense their private preference for the resource is. Clearly, such a mechanism may be advantageous when preference intensities/urgencies in the population are heterogeneous and not perfectly correlated over time. Indeed, (Elokda et al., 2023) show that farsighted Nash equilibria are almost fully efficient with respect to the private urgency of the users.
@@
To date, the literature on karma focused on the theoretical properties of the mechanism compared with random allocation and other domain-specific schemes such as dynamic pricing or max-min allocation (Vuppalapati et al., 2023; Elokda et al., 2022, 2023), all predicted on idealized assumptions such as rationality, far-sightedness and perfect adoption. In this paper, we complement the pre-existent literature by a first behavioral investigation with real humans who do not necessarily follow such idealized assumptions. We want to know whether humans find it natural to adapt the mechanism, and whether inexperienced players are able to realize efficiency gains. To this end, we conducted a controlled experiment to better understand whether and how effectively karma might be implemented in human populations as a resource governance mechanism. In our analysis we focus on the overall efficiency effects of karma as well as on distributional consequences and associated fairness properties of the mechanism. In our treatments, we vary the nature of the dynamic preference intensity distributions (more frequent and less intense versus less frequent and more intense) and the auction process of the karma mechanism (binary bidding versus full bidding). We use random allocation as a benchmark to compare our findings to. This domain-independent benchmark is representative of common schemes that are unaware of the individual private preferences, including fixed turn-taking schemes and schemes using fixed-value tokens.
Our main findings summarize as follows. Overall, compared with random allocation, karma leads to substantial efficiency gains, and these gains benefit almost all participants. In fact, only non-adopters, that is, individuals who do not participate in karma bidding actively themselves, do not benefit and may be worse off. The treatment variations suggest that benefits are particularly pronounced in situations when preference intensities are dynamically more intense and less frequent, and the bidding scheme is designed to be minimal (i.e., binary). These findings provide a first benchmark that karma may be used beneficially in human interactions. Our study also points in several directions for further theoretical and experimental investigation.
2. Experimental Methods
@@
2.1. The Game
The game we study is one that proceeds with $N$ participants over $T$-many rounds. All participants receive an initial endowment of karma $k(0)=k(1)={k_{\textup{init}}}\in\mathbb{N}$ and an initial game score $s(0)=0$. At each round $t\in\{1,\dots,T\}$, participants are randomly matched in pairs to compete over a shared resource. Each participant is given an urgency value $u(t)\in\{{u_{\textup{l}}},\>{u_{\textup{h}}}\}$ that is drawn randomly and independently from a process $\mathbb{P}(u)$ that is identical for all players, and must place a bid $b(t)\leq k(t)$. In each pairwise matching, the higher bidder gets allocated the resource: the score $s(t)=s(t-1)+u(t)$ increases by the urgency and the bid is collected as payment. The lower bidder does not get allocated the resource: the score $s(t)=s(t-1)$ does not increase and no payment is collected. Ties are settled randomly. To keep the total amount of karma constant, the total payment ${p_{\textup{tot}}}$ is uniformly redistributed to the participants in an integer preserving manner. In case ${p_{\textup{tot}}}$ is not divisible by $N$, $\left\lfloor\frac{{p_{\textup{tot}}}}{N}\right\rfloor$ is redistributed to everyone, and a random subset of ${p_{\textup{tot}}}-N\left\lfloor\frac{{p_{\textup{tot}}}}{N}\right\rfloor$ participants receive one additional karma. Moreover, each participant is allowed a maximum karma level ${k_{\textup{max}}}$ that the redistribution respects. Participants with ${k_{\textup{max}}}$ do not receive additional redistribution which instead gets issued uniformly to the others. After the redistribution the game proceeds to the next round $t+1$.
Table 3 in Appendix A summarizes the notation introduced above.
2.2. Treatments
We follow a 2x2 factorial treatment design, where we vary the dynamic urgency process of the participants and the richness of the karma scheme. Table 1 summarizes the treatment configurations schematically. For each of the four resulting treatments we run 5 independent game experiments with 20 participants per experimental game resulting in 100 subjects per treatment and 400 subjects in total.
Table 1. 2x2 experimental design.
@@ 2.2. Treatments
2.2.1. Urgency Process
The urgency processes in both treatment variations of low stake and high stake have the same low urgency ${u_{\textup{l}}}=1$, but differ in the magnitude and frequency of the high urgency (low stake: ${u_{\textup{h}}}=5$, ${\mathbb{P}}({u_{\textup{h}}})=0.5$; high stake: ${u_{\textup{h}}}=9$, ${\mathbb{P}}({u_{\textup{h}}})=0.25$). Both processes have the same urgency on average $\mathbb{E}(u)=3$, and therefore the same expected scores under random allocation. Low stake represents the case where the high urgency event is frequent but moderate, while in high stake the high urgency event is rare but severe.
2.2.2. Richness of Karma Scheme
In the treatment variation of binary, participants can only choose between two bid levels: 0 or $\left\lfloor\frac{k}{2}\right\rfloor$ (i.e., half their karma, rounded down to the nearest integer). In the treatment variation of full range, participants can choose any integer bid up to their karma. The binary treatment aims at investigating the behavioral effects of a reduced action space, whereby subjects either bid or not. Using tools from (Elokda et al., 2023), we designed the urgency processes in such a way that there do not result noteworthy efficiency differences associated with the Nash equilibria of the two schemes. Therefore, any differences observed in terms of outcomes between the two bidding schemes will be due to behavioral effects.
2.3. Experimental Implementation
The game is implemented as a real-time online experiment using oTree (Chen et al., 2016). Participants are given the opportunity to familiarize with the game over $5$ rounds that do no contribute to the game score, after which the main game proceeds with $T=50$ rounds. In each round, participants get drawn a random urgency. On the decision page, they must place a bid, and then a results page follows which provides feedback on the outcomes of that round. The bid must be placed within $10$ seconds, otherwise the participant is signalled as inactive, and the bid defaults to zero. The results page communicates the outcome of the round in terms of whether priority is granted, the karma payment and redistribution, and the updated karma balance and game score. In addition, it gives feedback on the opposing bid (but not the opposing urgency and karma). Figures 45 in Appendix A show examples of these pages.

### v2

# Dynamic Recourse Allocation with Karma: An Experimental Study

@@ Abstract
A system of non-tradable credits that flow between individuals like karma, hence proposed under that name, is a mechanism for repeated resource allocation that comes with attractive efficiency and fairness properties, in theory. In this study, we test karma in an online experiment in which human subjects repeatedly compete for a resource with time-varying and stochastic individual preferences or urgency to acquire the resource. We confirm that karma has significant and sustained welfare benefits even in a population with no prior training. We identify mechanism usage in contexts with sporadic high urgency, more so than with frequent moderate urgency, and implemented as a simple (binary) karma bidding scheme as particularly effective for welfare improvements: relatively larger aggregate efficiency gains are realized that are (almost) Pareto superior. These findings provide guidance for further testing and for future implementation plans of such mechanisms in the real world.
keywords:
Behavioral economics , Repeated allocation , Karma economy , Artificial currency
journal: Games and Economic Behavior
[…]
[ifa] organization=Automatic Contol Lab, ETH Zürich,addressline=Physikstrasse 3, city=Zürich, postcode=8092, country=Switzerland
[…]
[uzh] organization=Department of Sociology, University of Zürich,addressline=Andreastrasse 15, city=Zürich, postcode=8050, country=Switzerland
1 Introduction
@@
Figure 1: Dynamic resource allocation with karma involving three cars and a repeated sequence of three encounters. Start with the encounter in the top center. The high-urgency lila car has a current karma account of 9 and bids 6, thus outbidding the low-urgency blue car whose karma account is also 9 but bids 2. The bids of the lila car (and, in general, of other cars winning in parallel encounters) are paid and redistributed. As a result, the blue car’s karma account goes up by 3 to 12. Let’s move along clockwise. Blue now happens to have high urgency and bids 6, thus outbidding and getting priority over orange whose karma account is 9 and bids 3. Now the orange karma goes up by the redistribution share of 3 to 12. Let’s move along clockwise. Orange now has high urgency and bids 6, thus outbidding low-urgency lila who has 6 karma left and bids 1. Thus the circle closes, et cetera, et cetera.
Efficiency and fairness in determining who gets what and when are the two major objectives in resource allocation situations under scarcity, and many interesting mechanism and market design solutions have been proposed (Roth, 2015). An important class of allocation problems is when goods are repeatedly and indefinitely allocated amongst a fixed population: for example, farmers require daily access to shared groundwater resources, commuters require regular access to roads, students require frequent access to scarce computing clusters, food banks require daily access to food donations, etc. In situations like these, it is often the case that one person relative to another gets higher utility from the good in one period but lower utility in another, which we shall refer to as time-varying levels of urgency. A recent innovation to address exactly those kinds of allocation problems goes under the name karma (Vishnumurthy et al., 2003; Elokda et al., 2023; Vuppalapati et al., 2023). The karma mechanism mirrors what Western popular culture associates with the notions of karma and samsara stemming from Indian religions according to which one’s deeds in the present (karma) affect the quality of one’s future life (phala) and thereafter (samsara) (Reichenbach, 1990; Kyabgon, 2015). The mechanism proposed in that literature is implemented via individual accounts of non-tradable credits called karma. Individuals may bid some amount from their current account at every instance of resource allocation. In the baseline implementation of the mechanism, the highest bidders win and obtain priority for the resource, and must pay their bids, which are then redistributed in the population [^1] This corresponds to the first-price auction implementation of the mechanism. Other auctions, like second-price, have also been discussed.. Figure 1 illustrates how the karma mechanism works out to the benefit of everyone at hand of an example involving three infinitely re-occurring road intersection encounters.
The karma mechanism is simple and appealing, but not yet frequently used in practice. An exception is the ‘choice system’ proposed by Prendergast (2022), which is adopted by U.S. food banks, where karma tokens are called ‘shares.’ Food banks under the choice system bid shares to obtain priority for available food donations, pay their bids upon winning based on a first-price auction mechanism, and the total payment of shares is redistributed at the end of each day [^2] Indeed, Prendergast (2022) discusses alternatives to first-price auctions and their potential benefits, but mentions that the real-world implementation partners had a strong preference for first-price owed to its transparency and simplicity.. This real-world example is emblematic of the kind of resource allocations for which karma is suited: the use of monetary transfers is deemed inappropriate or highly undesirable; the allocations are repeated frequently (i.e., daily); and, importantly, there is no finite horizon in sight for when the allocations will cease to exist. Indeed, the design choice of redistributing the paid shares is critical to address the infinite repetition of the allocation: by forming a ‘closed economy’ in which total shares are preserved over time, a stationary regime can be reached and henceforth repeated indefinitely.
The idea of sacrificing resource consumption today in favor of future consumption in periods of higher urgency is intuitive, and it is not surprising that this same idea underpins several related lines of works, including linking decisions across periods (Jackson and Sonnenschein, 2007; Hortala-Vallve, 2010; Escobar and Toikka, 2013), trading favors (Möbius, 2001; Olszewski and Safronov, 2018b, a; Leo, 2017), token or artificial currency-based mechanisms (Johnson et al., 2014; Gorokh et al., 2021a, b; Banerjee et al., 2023), and trading votes (Casella and Macé, 2021; Casella, 2005; Casella et al., 2006; Hortala-Vallve and Llorente-Saguer, 2010; Hortala-Vallve, 2012; Casella and Palfrey, 2019, 2021; Macías, 2024). With respect to these works, which are discussed in Section 2, karma and the aforementioned choice system have the distinguishing feature of forming closed economies that are particularly suited for infinitely repeated resource allocations.
The choice system, which has indeed resulted in significant aggregate gains in terms of food bank participation and fluidity of donations (Prendergast, 2022), provides empirical evidence that karma mechanisms can be successful in practice. On the theory side, recent works have studied game-theoretical models of karma mechanisms in simplified, single repeated resource allocation settings (Elokda et al., 2023, 2024). These works have been motivated by the use of karma as a public policy instrument that is an equitable alternative to classical (monetary) congestion pricing policies, e.g., for allocating priority roads and other public infrastructures. However, in order to realize the potential benefits of karma-based policies, and employ these policies effectively in human populations, systematic behavioral evidence is needed. The present paper contributes to this emergent strand of literature a first controlled experimental test of karma. We conduct an experiment on karma with different bidding schemes and different urgency processes in order to better understand the behavioral uptake of karma by human actors and its efficiency and distributional consequences.
@@
The theoretical predictions of Elokda et al. (2023, 2024) are based on the solution concept of the Stationary Nash Equilibrium (SNE): a compact, time-invariant predictor of optimal rational behavior, which is guaranteed to exist due to the preservation of karma. Elokda et al. (2023) develop computation tools for the SNE and show that farsighted SNE are almost fully efficient with respect to the private urgency of the users. These predictions are based on idealized assumptions including rationality, far-sightedness, infinite population, perfect adoption, and that the stationary conditions for which the SNE is optimal are reached. In this paper, we perform a behavioral investigation with real humans who do not necessarily follow such idealized assumptions. We want to know whether humans find it natural to adopt the mechanism, whether inexperienced players are able to realize efficiency gains, and whether stationarity is attained in practice. In our analysis we focus on the overall efficiency effects of karma as well as on distributional consequences and associated fairness properties of the mechanism. In our treatments, we vary the nature of the dynamic urgency distributions (more frequent and less intense versus less frequent and more intense) and the auction process of the karma mechanism (binary bidding versus full bidding). We use random allocation as a benchmark to compare our findings to. This domain-independent benchmark is representative of common schemes that are unaware of the individual private urgency, including fixed turn-taking schemes and schemes using fixed-value tokens.
The main insights of our experiments summarize as follows:
  • 1.
    Almost all participants benefit in the karma allocation compared with random allocation: this is the case for 90% of the participants, and the remaining 10% are mostly non-adopters who do not participate in karma bidding actively themselves. If we look at active participants only, karma led to an almost Pareto improvement.
  • 2.
    The Pareto improvement occurs despite the relatively low level of training and commitment by the online participants.
  • 3.
    The realized benefits, while greater than in random allocation, fall short of theoretical Nash predictions. Analysis of the bidding behaviors reveals that the main deviation to Nash behavior takes the form of irrational over-bidding in low urgency rounds.
  • 4.
    Benefits are particularly pronounced in situations when preference intensities are dynamically more intense and less frequent, and the bidding scheme is designed to be minimal (i.e., binary).
  • 5.
    The realized variations in the karma and bid distributions over time are small and comparable in magnitude to the variations that would be attained if all participants followed stationary Nash behavior. This suggests that an approximately stationary regime is reached despite the presence of noisy bidding behaviors. Moreover, the variations are smaller under binary bidding than full bidding, suggesting that the karma auctions are particularly predictable under binary bidding.
These findings provide a first benchmark that karma may be used beneficially and robustly in human interactions. Our study also points in several directions for further theoretical and experimental investigation.
2 Related Mechanisms
In this section, we highlight how karma mechanisms differ from several previously proposed mechanisms that share the same intuitive idea of trading off between present or future access to resources.
Linking decisions
Mechanisms based on linking decisions (Jackson and Sonnenschein, 2007; Hortala-Vallve, 2010; Escobar and Toikka, 2013) rely on correlating each individual’s reports over time with publicly known distributions of the private urgency, and punishing those individuals that deviate. This requires to keep track of individual identities and histories, and we view karma and other tokens as memory-efficient instruments to link decisions that are capable of scaling in large populations.
Trading favors
Mechanisms based on trading favors (Möbius, 2001; Olszewski and Safronov, 2018b) rely on simple book-keeping of favors owed, but these classical mechanisms are tailored to truthful reporting of one’s availability to grant favors with no regard to time-varying private urgency. Leo (2017) addresses time-varying urgency specifically for two individuals taking turns to perform chores, while Olszewski and Safronov (2018a) addresses time-varying urgency in more general settings using probabilistic exchange of karma-like tokens called ‘chips’. However, Olszewski and Safronov (2018a)’s mechanism depends on the individual preference distributions in a complex manner and thus does not scale naturally.
Tokens and artificial currency
Karma-like instruments have been previously referred to as vouchers, tokens, scrips, or artificial currency. We distinguish between token-based mechanisms in which the value of the resource is fixed in tokens (Johnson et al., 2014) (typically one resource unit is worth one token); and artificial currency-based mechanisms in which, like karma, the value of the resource is determined in an auction-like mechanism (Gorokh et al., 2021a, b; Banerjee et al., 2023). Token-based mechanisms are not well suited to elicit time-varying private urgency; whereas most previously proposed artificial currency-based mechanisms are tailored to finite resource repetitions: individuals are issued an initial budget of currency to spend over the finite horizon (with no redistribution or other forms of currency exchange). The non-preservation of total system currency makes these mechanisms less well-suited to infinite resource repetitions: they would require a periodic central endowment of currency (e.g., every month or year), do not forgive mistakes leading to early depletion of currency, and, importantly, lead to non-stationary settings in which optimal strategies depend explicitly on the time left in the horizon.
Trading votes
Trading votes across issues or proposals is an intuitively appealing and practically prevalent practice, yet it remains unclear to what extent vote trading improves welfare and how to design vote trading mechanisms optimally, as pointed out in Casella and Macé (2021)’s recent review. Casella and Macé (2021) distinguish between two types of vote trading: those in which votes are traded with other voters (Casella and Palfrey, 2019, 2021); and those in which votes are traded individually with one’s future self (referred to as storable (Casella, 2005) or qualitative votes (Hortala-Vallve, 2012)). The latter type of storable votes, which yields particularly favorable efficiency gains in comparison to the other types of vote trading (Casella and Macé, 2021), is closely related to the aforementioned class of (finite-horizon) artificial currency mechanisms: voters are issued an initial budget of votes to cast in a (small) finite number of issues (Casella, 2005; Casella et al., 2006; Hortala-Vallve, 2012). One recently proposed mechanism by Macías (2024) resembles karma more closely: in this mechanism, votes are “paid” by the majority voters and subsequently redistributed, however, Macías (2024) studies a two voter only model. Our study thus complements the literature on storable votes, as we are motivated by resource allocations that are repeated more frequently and typically involve more players than in voting. In contrast to previous experimental studies on vote trading (Casella et al., 2006; Hortala-Vallve and Llorente-Saguer, 2010; Casella and Palfrey, 2019), our experiments involve significantly more rounds and larger groups.
3 Experimental Methods
@@
3.1 The Game
The game we study is one that proceeds with $N$ participants over $T$-many rounds. All participants receive an initial endowment of karma $k(0)=k(1)={k_{\textup{init}}}\in\mathbb{N}$ and an initial game score $s(0)=0$. At each round $t\in\{1,\dots,T\}$, participants are randomly matched in pairs to compete over a shared resource. Each participant is given an urgency value $u(t)\in\{{u_{\textup{l}}},\>{u_{\textup{h}}}\}$ that is drawn randomly and independently from a process $\mathbb{P}(u)$ that is identical for all players, and must place a bid $b(t)\leq k(t)$. In each pairwise matching, the higher bidder gets allocated the resource: the score $s(t)=s(t-1)+u(t)$ increases by the urgency and the bid is collected as payment. The lower bidder does not get allocated the resource: the score $s(t)=s(t-1)$ does not increase and no payment is collected. Ties are settled randomly. To keep the total amount of karma constant, at the end of each round, the total collected payment ${p_{\textup{tot}}}$ is uniformly redistributed to all participants in an integer preserving manner. In case ${p_{\textup{tot}}}$ is not divisible by $N$, $\left\lfloor\frac{{p_{\textup{tot}}}}{N}\right\rfloor$ is redistributed to everyone, and a random subset of ${p_{\textup{tot}}}-N\left\lfloor\frac{{p_{\textup{tot}}}}{N}\right\rfloor$ participants receive one additional karma. Moreover, each participant is allowed a maximum karma level ${k_{\textup{max}}}$ that the redistribution respects. Participants with ${k_{\textup{max}}}$ do not receive additional redistribution which instead gets issued uniformly to the others. After the redistribution the game proceeds to the next round $t+1$.
Table 3 in A summarizes the notation introduced above.
3.2 Treatments
We follow a two-by-two factorial treatment design, where we vary the dynamic urgency process of the participants and the richness of the karma scheme. Table 1 summarizes the treatment configurations schematically. For each of the four resulting treatments we run 5 independent game experiments with 20 participants per experimental game resulting in 100 subjects per treatment and 400 subjects in total.
Table 1: 2x2 experimental design.
@@ 3.2 Treatments
3.2.1 Urgency Process
The urgency processes in both treatment variations of low stake and high stake have the same low urgency ${u_{\textup{l}}}=1$, but differ in the magnitude and frequency of the high urgency (low stake: ${u_{\textup{h}}}=5$, ${\mathbb{P}}({u_{\textup{h}}})=0.5$; high stake: ${u_{\textup{h}}}=9$, ${\mathbb{P}}({u_{\textup{h}}})=0.25$). The motivation for these treatment variations is to investigate behavioral effects under urgency processes of different dynamic nature. Both processes have the same urgency on average $\mathbb{E}(u)=3$, and therefore the same expected scores under random allocation. Low stake represents the case where the high urgency event is frequent but moderate, while in high stake the high urgency event is rare but severe. Notice that due to the different dynamic nature of the processes, there is a greater potential to benefit over random allocation in the high stake process than in the low stake process (cf. Nash efficiency gains in Figure 2).
3.2.2 Richness of Karma Scheme
In the treatment variation of binary, participants can only choose between two bid levels: 00 or $\left\lfloor\frac{k}{2}\right\rfloor$ (i.e., half their karma, rounded down to the nearest integer). In the treatment variation of full range, participants can choose any integer bid up to their karma. The motivation of these treatment variations is to investigate the behavioral effects of a reduced action space, i.e., binary, where subjects either bid or not. The particular design choice to restrict bids to 00 or $\left\lfloor\frac{k}{2}\right\rfloor$ in the binary scheme was guided by theoretical Nash predictions: using tools from Elokda et al. (2023), these binary bid levels were predicted to achieve almost the same efficiency at the Nash equilibrium as in full range bidding, for the urgency processes considered (cf. Nash efficiency gains in Figure 2). Therefore, the simpler binary scheme does not trade-off performance, in theory, and any differences observed in the outcomes of the two bidding schemes will be due to behavioral effects.
3.3 Experimental Implementation
The game is implemented as a real-time online experiment using oTree (Chen et al., 2016). Participants are given the opportunity to familiarize with the game over $5$ rounds that do no contribute to the game score, after which the main game proceeds with $T=50$ rounds. In each round, participants get drawn a random urgency. On the decision page, they must place a bid, and then a results page follows which provides feedback on the outcomes of that round. The bid must be placed within $10$ seconds, otherwise the participant is signalled as inactive, and the bid defaults to zero. The results page communicates the outcome of the round in terms of whether priority is granted, the karma payment and redistribution, and the updated karma balance and game score. In addition, it gives feedback on the opposing bid (but not the opposing urgency and karma). Figures 78 in A show examples of these pages.